OpenStudy (idealist10):

How to integrate ln abs(1+x) dx?

OpenStudy (idealist10):

@campbell_st @e.mccormick @Destinymasha @sammixboo

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\ln \left| 1+x \right| dx ?\]

OpenStudy (anonymous):

like this?

OpenStudy (freckles):

you can use integration by parts

OpenStudy (freckles):

and I see from a previous question you are doing it from x=0 to x=1 so we have to not worry about | |

OpenStudy (freckles):

oops this question is closed